### v1

@@ 7.5 Continuous Testing and Delivery of Quantum Services
With an adoption of agile software engineering in quantum software development context [66], there is a need for light and adaptive methods to ensure a continuous development and delivery of quantum software services. The literature suggested a lack of solutions on testing the software services. The continuous testing and continuous delivery can help CT/CD to test the services against QSRs more effectively and deliver them rapidly. Quantum service testing can involve simulation or regression tests against the QSRs. Quantum service testing involves various approaches, including simulation and regression tests against QSRs. Given the unique characteristics of quantum computing, such as superposition and entanglement, traditional testing methods may not suffice. Therefore, specialised testing strategies are essential to ensure the reliability and performance of quantum services.
Furthermore, integrating CT/CD practices into the quantum software development lifecycle facilitates early detection of issues, reduces feedback times, improves visibility of product quality, and accelerates time-to-market [68]. This integration is crucial for maintaining the agility and adaptability required in the rapidly evolving field of quantum computing
@@
Continuous Testing and Continuous Delivery (CT/CD) relies on agile software engineering methods to help deliver quantum software services rapidly and reliably. CT/CD can provide strategic benefits to vendors by adding new services to their quantum platforms.
@@
Large Language Models and Quantum Service Engineering: Although we did not find any explicit evidence, specifically published research on synergising generative artificial intelligence (including the large language models (LLMs)) and quantum service development. However, some recent research studies indicate a growing trend on exploiting LLMs to design and develop quantum software. Recent work shows that LLMs can mimic the behaviour of quantum circuits and predict how quantum states evolve, thus enabling a faster and resource efficient alternative to traditional quantum simulators [69]. Purpose-built LLMs such as the GroverGPT, have claimed to achieve near-perfect accuracy when replicating Grover’s search algorithm on systems with up to 20 qubits [70]. Beyond the general purpose simulations, LLM-based solution are improving the quality of quantum source code through tools such as LintQ-LLM, which can detect, locate, and explain programming errors in Qiskit more effectively than standard static analysis methods [71]. They are also being used to help developers keep up with rapid changes in quantum programming libraries—such as migrating between Qiskit versions—by automatically identifying and classifying common refactoring scenarios [72]. The collective impact of these advancements indicates that LLMs, already recognised for their role in software engineering and architecting [73], are likely to become central to quantum software engineering—enabling code generation and testing, as well as supporting the simulation, development, and maintenance of quantum software services.
@@
QCaaS in the era of LLMs: Recent studies show LLMs can simulate quantum circuits, enhance code quality, and support library migration in quantum software development. While direct research on combining LLMs with quantum service computing is scarce, emerging tools like GroverGPT and LintQ-LLM indicate LLMs may soon become central to quantum software engineering.
8 Related Survey-based and Empirical Studies
@@
Interpretation of the Comparative Analysis: For an objective comparison, Table 3 presents each study using five-point self-explanatory criteria including (a) type of study (adopted from [21]), (b) focus of study, (c) core findings, (d) QSE activity supported by the study, and (e) year of publication, each exemplified below. For example, the study [19] presents a systematic review, conducted in 2021, as part of evidence-based software engineering to focus on architecting quantum software. The SLR presents the core findings about architectural life cycle and state-of-the-art on architectural modeling, patterns, and tools to architect quantum software. The SLR focuses on design and architecting activities of QSE life cycle and indicates the needs for future research on quantum-specific professional expertise (human roles in QSE) to develop quantum software.
Table 3: A Review of the Relevant Secondary Studies
@@ 8.1 Quantum Software Engineering
Quantum Software Engineering (QSE) has emerged as the most recent genre of software engineering (SE) that allows practitioners to adopt a process-centric and systematic approach to develop software-intensive systems and applications for QC [35, 75]. We have found a total of 5 survey-based studies, summarised in Table 3 to compare the scope and contributions of our proposed research. Some recently published survey-based studies on QSE highlight that existing SE principles and practices can be tailored to develop quantum software, however, issues specific to QC, such as operationalising QuBits/QuGates, quantum domain engineering, and quantum-classic software split, require new engineering methods to tackle these challenges [75]. To derive new methods and processes, the QSE research community is striving to organize a body of knowledge and academic collaborations that can be stimulated via dedicated publication forums to disseminate quantum software specific research including, but not limited to IEEE QSW  [^3] IEEE International Conference on Quantum Software. https://services.conferences.computer.org/2025/qsw/qsw-call-for-papers/ and ACM TOSEM’s Open Continuous Special Section on Quantum SE [^4] ACM Transactions on Software Engineering and Methodology. https://dl.acm.org/journal/tosem/quantumse . Academic discussions at the publication fora and published results [1][4] have highlighted that in addition to the engineering principle and practices, human knowledge and expertise require fundamental knowledge of quantum mechanics to effectively design algorithmic solutions for QSE projects [25]. The current generation of software architects and developers who lack the foundational knowledge of quantum mechanics may be hindered or find themselves under-prepared for quantum software development [19]. Software architecture is being viewed as a solution that can abstract complexities of implementation by representing software to be developed as architectural components and connectors. A recently conducted systematic review of architectural solutions puts forward five architecting activities to guide software designers to engineer quantum software solutions in an incremental manner [19]. Architectural solution(s) can also empower quantum software engineers to exploit architectural knowledge in terms of patterns (i.e., reuse knowledge), modelling notations (i.e., software representation), and tools (i.e., automation) to address the challenges of QSE. In addition to architecting, there is a need for empirically grounded guidelines and solutions to address the aspects of quantum software programming, model-driven development, and quality assurance in QSE [74].
@@
Despite the recent advancements in Quantum Software Engineering (QSE) discipline, a multitude of technical and human skills related challenges persist that require innovative solutions and empirical research to address them. The study in [76] presents the results of a systematic mapping study of the current state of QSE research and identifies trends such as the most investigated topics, types and number of studies, main reported results, and the most studied quantum computing tools and frameworks. The study concludes by exploring the research community’s interest in QSE, how the discipline is evolving, and implications of research results are as a centralised source of information to facilitate knowledge transfer for researchers and practitioners. The authors of [81] have conducted a two-phase empirical study to investigate the impcat of qunatum programming language on qunatum software development. Specifically, at the first stage, the authors mined data from GitHub repositories to investigate the quantum programming frameworks used by developers for qunatum software development. Secondly, the authors conducted a survey targeting the contributors of these repositories to capture their perspectives on the current adoption and challenges of quantum programming. The study [20] conducted in-depth semi-structured interviews to investigate practationers’ view on the issues and challenges of quantum software engineering. The insights from interviews provide developers’ perspectives such as the motivations, challenges, and outlooks within the rapidly evolving field of QSE. Despite promising advancements in QSE, significant challenges persist and demand attention. De Stefano et al. in [76] analysed widely-used quantum tools, offering valuable insights into their capabilities and limitations. Arias et al. in [18] explored security issues in quantum programming languages and hardware, emphasising their impact on software development. The study highlighted critical vulnerabilities and advocated for the early integration of security measures to mitigate potential exploitation. Murillo et al. [77] further emphasised the need for robust frameworks and comprehensive security protocols. As quantum computing continues to advance, integrating QSE into practical applications becomes increasingly critical, requiring a thorough understanding of both its potential and associated risks.
@@ 8.2 Quantum Service Computing
Quantum service-orientation is a recent trend initially pushed by QC vendors to allow developers who can compose and invoke software services on remotely hosted quantum systems and infrastructures [16]. We identified a total of 2 survey-based studies on quantum service computing highlighted in Table 3. Quantum service computing initiatives, such as Amazon Braket and Q experience, have paved the way for academic research to propose solutions for quantum cloud, quantum services computing, and quantum as a service. Recent research reviews inform about the current progress and emerging challenges for quantum services computing that include but are not limited to hardware availability, quantum noise, and quantum-classic split [80]. Software researchers are striving to synergise existing methods of microservicing and quantum software development, such as Quantum DevOps to develop quantum microservices [66].
Researchers on quantum cloud computing and distributed quantum computing highlighted both significant advancements and ongoing challenges in leveraging quantum technologies for practical use. Both fields emphasised the importance of integrating quantum computing with existing infrastructures to expand computational capabilities and overcome current limitations. The study in [16] provided a comprehensive review of quantum cloud computing, examining models, platforms, technologies, and critical issues such as resource management, security, and privacy. Their study identified open problems and outlined opportunities for future exploration. Similarly, the study [80] explored the challenges and unresolved issues in distributed quantum computing, offering insights from a computing and communications perspective. In addition, researchers on quantum cloud computing (QCC) and distributed quantum computing highlighted both significant advancements and ongoing challenges in leveraging quantum technologies for practical use. Both fields emphasised the importance of integrating quantum computing with existing infrastructures to expand computational capabilities and overcome current limitations. Nguyen et al. [78] provided a comprehensive review of quantum cloud computing, examining models, platforms, technologies, and critical issues such as resource management, security, and privacy. Their study identified open problems and outlined opportunities for future exploration. Similarly, Caleffi et al. []explored the challenges and unresolved issues in distributed quantum computing, offering insights from a computing and communications perspective.
@@ 9.1.1 Implications for Research
The results of this study are presented as structured catalogues (Table 2) and illustrative visualisations (Figure 2 - Figure 7) that provide a synthesized knowledge on research state-of-the-art and vision for future needed research. Core findings of the study organised as service lifecycle activities in Section 5 and their integration into a reference architecture in Section 6 can inform researchers about leveraging existing SOA principles, practices, and patterns in the context of emerging design challenges of QCaaS. Specifically, the reference architecture derived from existing literature provides a blue-print that harnessesăpatterns and tactics from classical software engineering (layered architecture, service facade) that can be integrated with quantum-specificăpatterns such as the quantum API gateway and quantum-classic split. An illustrative discussion of emerging trends can provide a roadmap as in Figure 9 for future research involving empiricism such as surveys, interviews, or repository analyses research to develop evidence-based taxonomies to architect and implement QCaaS. The study results also indicate that the ‘Manage phase of the SOA lifecycle [26] is currently underexplored, offering potential for future research on quantum-specific service monitoring, governance, and quality management. Researchers could explore how LLMs might automate tasks such as generating quantum service models from natural language, completing quantum code, migrating code between frameworks, or simulating quantum circuits - speeding up the development via human-bot collaborative software engineering for QCaaS.
@@ 9.1.2 Implications for Practice
The results of this study provide evidence-based guidelines for practitioners, including quantum software architects and developers about the QSRs that guide quantum software and service development process (Section 8). The pattern catalogue and thumbnails in Figure 6 represent a concentrated wisdom of quantum service that enables software designers and architects to rely on reuse knowledge and best practices to engineer QCaaS. The study indicates that QSE requires dedicate industrial roles such as qunatum domain engineer, quantum simulation manager, quantum algorithm desinger in an industrial scale development of unatum software as in Figure 9. The predominance of Python-based programming frameworks and qunatum service execution environments such as Qiskit and Amazon Braket reflects the gradual maturity quantum software/service ecosystem, reduces adoption risk, and enables practitioners to leverage established guidelines and concrete frameworks for service assembly and deployment. The study can help practitioners with the structured knowledge, patterns, and architectural guidance needed to build reliable and efficient quantum services, helping to enable and enhance software engineering for QCaaS.
@@ 9.2.1 Threats to Internal Validity
Internal validity addresses whether the results of the study (e.g., service lifecycle phases, and derived reference architecture) are influenced only by the variables under study (e.g., SMS process, ABD method). Potential threats to internal validity may arise from factors including, but are not limited to literature selection bias, the subjectivity of quality assessment criteria, and inconsistencies in data extraction that have been presented in Section 3. We tried to minimise these threats by following a structured process (Figure 3), specifically customising the search string from Listing 1 as per individual EDS by following the guidelines in [41]. Moreover, to ensure a multistep quality assessment of each individual study, we derived a qulaity assessment checklist, as presented in Table 1. Finally, to minimise the threat of omitting key information from the literature, we created a structured template as the data extraction form Appendix A - Table 4 representing 11 data points to extract and record data from each primary study. The data extraction template provided a foundation to synthesise the data and document the study results.
@@ 9.2.2 Threats to External Validity
External validity concerns the extent to which the study’s findings can be generalised beyond the immediate scope of this research, including to different quantum platforms, service models, or development environments. These threats can include the dominance of academic research over industrial perspectives/solutions, temporal restrictions of the reviewed literature, and the narrow evidence base, i.e., 41 studies as the underlying evidence for the proposed reference architecture. As the findings are derived primarily from peer-reviewed academic sources (Appendix A - Table LABEL:tab:study-list), they lack practitioners’ perspective and industry-based insights for quantum service architecting and development. One possible way to minimise this threat is by using a mixed-method approach [83] that uses findings from literature to derive an architecture and develops a proof-of-the-concept validation as in Figure 8. The limitation can be further mitigated with future work that relies on survey research with quantum service developers and software engineers to validate and complement the findings of the study. Furthermore, extending the evidence base through approaches such as repository mining or practitioner-oriented evaluation of the reference architecture can provide stronger empirical grounding and enhance the generalisability of the proposed architecture.
@@ 9.2.3 Threats to Construct Validity
Construct validity addresses the extent to which a study accurately measures what it is intended to measure, ensuring that the research questions, methodological details, etc. are clearly defined and appropriately operationalised. In the context of this study, construct validity relates to whether the selected research questions, study selection criteria, and data extraction items (in Section 3) adequately capture the essential aspects of Quantum Computing as a Service (QCaaS) and its software engineering underpinnings. A key threat arises from the operationalisation of core constructs such as quantum service lifecycle, reference architecture, and quantum significant requirements (QSRs). Although these constructs have been defined based on established software engineering and quantum computing literature [35] [19], the rapid evolution of terminology within the quantum software engineering domain introduces a risk of conceptual drift, potentially limiting the generalisability of our findings to future research contexts.

### v2

@@ 7.5 Continuous Testing and Delivery of Services
With the adoption of agile methods in quantum software development [69], there is a need for light and adaptive approaches to enable continuous development and delivery of quantum services. The literature highlights a gap in testing quantum services. Continuous testing and delivery (CT/CD) can help validate services against QSRs more effectively and accelerate their deployment. Due to quantum-specific characteristics like superposition and entanglement, traditional testing methods are often insufficient, making specialised testing strategies essential to ensure the reliability and performance of quantum services. Furthermore, integrating CT/CD practices into the quantum software development lifecycle facilitates early detection of issues, reduces feedback times, improves contonuous delivery of QC services [71].
@@
Continuous Testing and Delivery (CT/CD) grounded in agile methods, enables the rapid and reliable delivery of quantum software services. It also offers strategic benefits to vendors by facilitating the addition of new services to their quantum platforms.
@@
Large Language Models and Quantum Service Engineering: Although we did not find any explicit evidence, specifically published research on synergising generative artificial intelligence (including the large language models (LLMs)) and quantum service development. However, some recent research studies indicate a growing trend on exploiting LLMs to design and develop quantum software. Recent work shows that LLMs can mimic the behaviour of quantum circuits and predict how quantum states evolve, thus enabling a faster and resource efficient alternative to traditional quantum simulators [72]. Purpose-built LLMs such as the GroverGPT, have claimed to achieve near-perfect accuracy when replicating Grover’s search algorithm on systems with up to 20 qubits [73]. Beyond the general purpose simulations, LLM-based solution are improving the quality of quantum source code through tools such as LintQ-LLM, which can detect, locate, and explain programming errors in Qiskit more effectively than standard static analysis methods [74]. They are also being used to help developers keep up with rapid changes in quantum programming libraries, such as migrating between Qiskit versions, by automatically identifying and classifying common refactoring scenarios [75]. The collective impact of these advancements indicates that LLMs, already recognised for their role in software engineering and architecting [76], are likely to become central to quantum software engineering - enabling code generation and testing, as well as supporting the simulation, development, and maintenance of quantum software services.
@@
LLMs for Quantum Service Engineering: Recent studies show LLMs can simulate quantum circuits, enhance code quality, and support library migration in quantum software development. While direct research on combining LLMs with quantum service computing is scarce, emerging tools like GroverGPT and LintQ-LLM indicate LLMs may soon become central to QSE.
8 Related Surveys and Empirical Studies
@@
Interpretation of the Comparative Analysis: For an objective comparison, Table 3 presents each study using five-point self-explanatory criteria including (a) type of study (adopted from [23]), (b) focus of study, (c) core findings, (d) QSE activity supported by the study, and (e) year of publication, each exemplified below. For example, the study [19] presents a systematic review, conducted in 2021, as part of evidence-based software engineering to focus on architecting quantum software. The SLR presents the core findings about architectural life cycle and state-of-the-art on architectural modeling, patterns, and tools to architect quantum software. The SLR focuses on design and architecting activities of QSE life cycle and indicates the needs for future research on quantum-specific professional expertise, i.e., human roles in QSE.
Table 3: Summary of the Core Findings of Most Relevant Secondary Studies
@@ 8.1 Quantum Software Engineering
QSE represents a recent and quantum-specific genre of the classical software engineering, providing practitioners with a systematic, process-oriented approach to developing software-intensive systems and applications for quantum computing [37] [78]. Our review identified eight survey-based studies, summarised in Table 3, which help compare the scope and contributions of our proposed research. Recent surveys indicate that while traditional SE principles can be adapted for quantum software development, QC-specific challenges including, but not limited to operationalising QuBits/QuGates and quantum domain engineering, remain largely unaddressed. To derive new methods and processes, the QSE research community is striving to organise a body of knowledge and academic collaborations that can be stimulated via dedicated publication forums to disseminate quantum software specific research including, but not limited to IEEE QSW  [^3] IEEE International Conference on Quantum Software. https://services.conferences.computer.org/2025/qsw/qsw-call-for-papers/ and ACM TOSEM’s Open Continuous Special Section on Quantum SE [^4] ACM Transactions on Software Engineering and Methodology. https://dl.acm.org/journal/tosem/quantumse. Academic discussions and published studies [1][4] highlight that beyond standard engineering principles, effective QSE requires a solid understanding of quantum mechanics [27]. Many current software architects and developers, lacking this foundational knowledge, may be under-prepared for quantum software development [19]. Software architecture offers a way to abstract implementation complexities by representing systems as components and connectors. A recent systematic review identifies five key architecting activities to guide designers in incrementally engineering quantum software solutions [19]. Architectural solution(s) can also empower quantum software engineers to exploit architectural knowledge in terms of patterns (i.e., reuse knowledge), modeling notations (i.e., software representation), and tools (i.e., automation) to address the challenges of QSE. In addition to architecting, there is a need for empirically grounded guidelines and solutions to address the aspects of quantum software programming, model-driven development, and quality assurance in QSE [77].
@@
Despite the recent advancements in QSE discipline, a multitude of technical and human skills related challenges persist that require innovative solutions and empirical research to address them. The study in [79] presents the results of a systematic mapping study of the current state of QSE research and identifies trends such as the most investigated topics, types and number of studies, main reported results, and the most studied quantum computing tools and frameworks. The study concludes by exploring the research community’s interest in QSE by exploring how the discipline is evolving, and what are the implications of research results as a source of information to facilitate knowledge transfer for researchers and practitioners. The authors of [84] have conducted a two-phase empirical study to investigate the impcat of qunatum programming language on qunatum software development. Specifically, at the first stage, the authors mined data from GitHub repositories to investigate the quantum programming frameworks used by developers for qunatum software development. Secondly, the authors conducted a survey targeting the contributors of these repositories to capture their perspectives on the current adoption and challenges of quantum programming. The study [20] conducted in-depth semi-structured interviews to investigate practationers’ view on the issues and challenges of quantum software engineering. The insights from interviews provide developers’ perspectives such as the motivations, challenges, and outlooks within the rapidly evolving field of QSE. Despite promising advancements in QSE, significant challenges persist and demand attention for research and development. De Stefano et al. in [79] analysed widely-used quantum tools, offering valuable insights into their capabilities and limitations. Arias et al. in [18] explored security issues in quantum programming languages and hardware, emphasising their impact on software development. The study highlighted critical vulnerabilities and advocated for the early integration of security measures to mitigate potential exploitation. Murillo et al. [80] further emphasised the need for robust frameworks and comprehensive security protocols.
@@ 8.2 Quantum Service Computing
Quantum service-orientation is an emerging trend driven by QC vendors, enabling developers to compose and invoke services on remotely hosted quantum systems [16]. Table 3 highlights four survey-based studies on quantum service computing. Platforms like Amazon Braket and IBM Q Experience have supported research on quantum cloud, quantum services, and QCaaS. Recent reviews highlight some prominent ongoing challenges, including limited hardware availability, quantum noise, and the quantum-classic split [83]. Researchers are exploring ways to integrate microservices and quantum software practices, such as Quantum DevOps, to support the development of quantum microservices [69].
Researchers on quantum cloud computing and distributed quantum computing highlighted both significant advancements and ongoing challenges in leveraging quantum technologies for practical use. Both fields emphasised the importance of integrating quantum computing with existing infrastructures to expand computational capabilities and overcome current limitations. In this context, the study reported in [16] provided a comprehensive review of quantum cloud computing, examining models, platforms, technologies, and critical issues such as resource management, security, and privacy. Their study identified open problems and outlined opportunities for future exploration. Similarly, the study [83] explored the challenges and unresolved issues in distributed quantum computing, offering insights from a computing and communications perspective. Nguyen et al. [81] provided a comprehensive review of quantum cloud computing, examining models, platforms, technologies, and critical issues such as resource management, security, and privacy. Their study identified open problems and outlined opportunities for future exploration. Similarly, Caleffi et al. [82] explored the challenges and unresolved issues in distributed quantum computing, offering insights from a computing and communications perspective.
@@ 9.1.1 Implications for Research
The results of this study are presented as structured catalogues (Table 2) and illustrative visualisations (Figure 2 - Figure 7) that provide a synthesised knowledge on research state-of-the-art and vision for future needed research. Core findings of the study organised as service lifecycle activities in Section 5 and their integration into a reference architecture in Section 6 can inform researchers about leveraging existing SOA principles, practices, and patterns in the context of emerging design challenges of QCaaS. Specifically, the reference architecture derived from existing literature provides a blue-print that harnessesăpatterns and tactics from classical software engineering (layered architecture, service facade) that can be integrated with quantum-specificăpatterns such as the quantum API gateway and quantum-classic split. An illustrative discussion of emerging trends can provide a roadmap as in Figure 9 for future research involving empiricism such as surveys, interviews, or repository analyses research to develop evidence-based taxonomies to architect and implement QCaaS. The study results also indicate that the ‘Manage’ phase of the SOA lifecycle [28] is currently underexplored, offering potential for future research on quantum-specific service monitoring, governance, and quality management. Researchers could explore how LLMs might automate tasks such as generating quantum service models from natural language, completing quantum code, migrating code between frameworks, or simulating quantum circuits - speeding up the development via human-bot collaborative software engineering for QCaaS.
@@ 9.1.2 Implications for Practice
The results of this study provide evidence-based guidelines for practitioners, including quantum software architects and developers about the QSRs that guide quantum software and service development process (Section 8). The pattern catalogue and thumbnails in Figure 6 represent a concentrated wisdom of quantum service that enables software designers and architects to rely on reuse knowledge and best practices to engineer QCaaS. The study indicates that QSE requires dedicated professional roles and expertise such as quantum domain engineer, quantum simulation manager, quantum algorithm desinger in an industrial scale development of unatum software as in Figure 9. The predominance of Python-based programming frameworks and qunatum service execution environments such as Qiskit and Amazon Braket reflects the gradual maturity quantum software/service ecosystem, reduces adoption risk, and enables practitioners to leverage established guidelines and concrete frameworks for service assembly and deployment. The study can help practitioners with the structured knowledge, patterns, and architectural blueprint needed to develop reliable and efficient quantum services for enabling and enhancing software engineering for QCaaS [76].
Figure 10: Synergising Quantum and Service Computing
@@ 9.2.1 Threats to Internal Validity
Internal validity addresses whether the results of the study (e.g., service lifecycle phases, and derived reference architecture) are influenced only by the variables under study (e.g., SMS process, ABD method). Potential threats to internal validity may arise from factors including, but are not limited to literature selection bias, the subjectivity of quality assessment criteria, and inconsistencies in data extraction that have been presented in Section 3. We tried to minimise these threats by following a structured process (Figure 3), specifically customising the search string from Listing 1 as per individual EDS by following the guidelines in [42]. Moreover, to ensure a multistep quality assessment of each individual study, we derived a qulaity assessment checklist, as presented in Table 1. Finally, to minimise the threat of omitting key information from the literature, we created a structured template as the data extraction form Appendix A - Table 4 representing 11 data points to extract and record data from each primary study.
@@ 9.2.2 Threats to External Validity
External validity concerns the extent to which the study’s findings can be generalised beyond the immediate scope of this research, including different QC platforms, service models, or development environments. These threats can include the dominance of academic research over industrial perspectives/solutions, temporal restrictions of the reviewed literature, and the narrow evidence base, i.e., 41 studies as the underlying evidence for the proposed reference architecture. As the findings are derived primarily from peer-reviewed academic sources (Appendix A - Table LABEL:tab:study-list), they lack practitioners’ perspective and industry-based insights for quantum service architecting and development. One possible way to minimise this threat is by using a mixed-method approach [86] that uses findings from literature to derive an architecture and develops a proof-of-the-concept validation as in Figure 8. The limitation can be further mitigated with future work that relies on survey research with quantum service developers and software engineers to validate and complement the findings of the study. Furthermore, extending the evidence base through approaches such as repository mining or practitioner-oriented evaluation of the reference architecture can provide stronger empirical grounding and enhance the generalisability of the proposed architecture.
@@ 9.2.3 Threats to Construct Validity
Construct validity addresses the extent to which a study accurately measures what it is intended to measure, ensuring that the research questions, methodological details, etc. are clearly defined and appropriately operationalised. In the context of this study, construct validity relates to whether the selected research questions, study selection criteria, and data extraction items (in Section 3) adequately capture the essential aspects of QCaaS and its software engineering foundations. A key threat arises from the operationalisation of core constructs such as quantum service lifecycle, reference architecture, and QSRs. Although these constructs have been defined based on established software engineering and quantum computing literature [37] [19], the rapid evolution of terminology within the quantum software engineering domain introduces a risk of conceptual drift, potentially limiting the generalisability of our findings to future research contexts.